# Parameter-Efficient Augment Plugin for Class-Incremental Learning

### v1

@@ 4.1 LoRA for Convolutional Layers
where $\mathbf{A}\in\mathbb{R}^{r\times C_{\text{in}}\times K\times K}$ projects the feature to a low-rank space. Then followed by kernel B:
@@
where $B\in\mathbb{R}^{C_{\text{out}}\times r\times 1\times 1}$ restores channel dimensionality. In summary, the forward pass with ConvLoRA is:
$\boldsymbol{h}^{\prime}=W_{\text{conv}}\ast\boldsymbol{x}+\frac{\alpha}{r}(B\ast(A\ast\boldsymbol{x})),$ (7)
where $\alpha$ is a scaling hyperparameter, $r$ is the low-rank.
4.2 Deploy LoRA as Task-Specific Plugins
LoRA cannot be directly applied in non-PTM settings, not only because PTMs provide diverse knowledge. Without a frozen PTM backbone, the output of the model changes with parameter updates, leading to unstable layer inputs for LoRA. However, LoRA becomes applicable in replay- and distillation-based methods, as these approaches effectively constrain the variation of the model’s outputs on old tasks, thereby ensuring that the input features to LoRA remain close to the distribution encountered during its training phase. Specifically, after training on task $t$, the output deviation of the model at any layer $\ell$ will not exceed $K_{\ell}\Gamma_{t}$ (The detailed proof is provided in the supplementary material). Here, $K_{\ell}$ is determined by hyperparameters such as the network structure and distillation temperature, while $\Gamma_{t}$ characterizes the intensity of distribution discrepancy induced by selected distillation and replay strategies. This upper bound limits the maximum feature drift magnitude of the model on the old tasks’ distribution.
Similar to expansion-based methods, our approach incorporates independent LoRA plugins for each task. Specifically, for each task $t$, we create a dedicated set of $k$ LoRA plugins for $\phi(\boldsymbol{x})$, denoted collectively as $L_{t}$. The $\phi(\boldsymbol{x})$ equipped with $L_{t}$ is denoted by $\phi^{L_{t}}(\boldsymbol{x})$. In DLC, LoRA plugins are deployed in the final layers of $\phi(\boldsymbol{x})$. Taking a CNN-based architecture as an example, the first LoRA plugin in $L_{t}$ is applied to the last convolutional layer of the $\phi(\boldsymbol{x})$. The remaining LoRA plugins are then sequentially deployed in a backward manner toward earlier layers. This design is motivated by two key considerations. First, features extracted from deeper layers are more semantically meaningful and are directly utilized for final classification decisions. Injecting task-specific residual adaptations via LoRA at this level allows for more direct enhancement of inter-class discriminability. Second, it demonstrates better compatibility with distillation-based objectives. When applied to shallow layers, the residual modifications introduced by LoRA would cause alterations in low-level features that propagate through the entire network, consequently amplifying the divergence in the final output logits. This increased discrepancy substantially raises the optimization difficulty for distillation losses, which rely on logit alignment between old and new models.
In DLC, the LoRA plugins and $\phi(\boldsymbol{x})$ undergo a decoupled, two-phase training procedure. Specifically, upon the arrival of a new task $t$, we first train only the $\phi(\boldsymbol{x})$ and classifier $\boldsymbol{W}$, keeping all LoRA plugins frozen, to facilitate adaptation to the new task distribution. After this phase, $\phi(\boldsymbol{x})$ is frozen to prevent knowledge drift, and the $L_{t}$ are trained with the same data. Upon completion, $L_{t}$ is permanently frozen to safeguard the acquired task-specific knowledge from forgetting. The training objective for all LoRA plugins incorporates the cross-entropy loss and the auxiliary loss [yan2021dynamically, zhou2022model] commonly employed in expansion-based methods, formulated as follows:
$\mathcal{L}_{\text{LoRA}}=\mathcal{L}_{\text{CE}}+\mathcal{L}_{\text{aux}}.$ (8)
The two-phase training procedure ensures that the LoRA plugins are trained without interference from distillation losses, enabling specialized learning for the current task and thereby improving its accuracy. Since the $\phi(\boldsymbol{x})$ remain frozen during LoRA updates, the training process of LoRA plugins exhibits stability. Furthermore, it prevents the $\phi(\boldsymbol{x})$’s training from being affected by LoRA, thereby maintaining the plug-and-play characteristic of DLC.
During the inference phase, all LoRA plugins are activated to extract their respective feature representations. These representations are then concatenated to form an integrated feature vector, which enhances the base model through a multi-task feature ensemble. For a linear classifier $\boldsymbol{W}$, the classification process can be formulated as:
@@ 4.2 Deploy LoRA as Task-Specific Plugins
where $T$ denotes the maximum index of known tasks. Compared to the model expansion, LoRA plugin expansion can bring significant parameter advantages. Specifically, for a convolutional layer ($C_{in}$, $C_{out}$, $K\times K$), the space complexity of its parameters is $\Theta(C_{in}C_{out}K^{2})$, while a ConvLoRA only occupies $\Theta(r(C_{in}K^{2}+C_{out}))$. For expanding one layer of CNN over $T$ tasks, considering that in a CNN in CIL, $C_{out}$ is generally greater than or equal to $C_{in}$, we set $C_{in}=C_{out}=C$, and $K$ can be regarded as a constant (usually 3 or 7). Since $r$ is low-rank, it can also be regarded as a constant (generally set to 8 or 16). Then, using LoRA expansion reduces the parameter space complexity of expanding one layer of CNN from $\Theta(TC^{2})$ to $\Theta(TC)$. Considering that $C$ is generally between 64 and 512, reducing its order of magnitude from quadratic to linear can bring a huge reduction in parameters.
@@ 4.3 Dynamic Weighting Unit
For a sample from task $t$, the LoRA plugins trained on other tasks is considered non-target plugins. Given that the replay buffer contains only a minimal fraction of the original data for each task, these non-target plugins are not trained on the full data from task $t$. In particular, plugins trained on tasks 1 to t1t-1 never encounter samples from the task $t$. Consequently, the residuals extracted by these plugins could be largely irrelevant to the task $t$, thus introducing uninformative features or even detrimental noise to the classifier.
@@
where the $\mathbf{\omega}_{\text{ideal}}$ is the ideal weight output, it has $\mathbf{\omega}^{(i)}_{\text{ideal-pos}}=1$ and $\mathbf{\omega}^{(i)}_{\text{ideal-pre}}=0$. In the ideal training case, the model has $\mathbf{\omega}^{(i)}_{\text{pos}}\to 1$ and $\mathbf{\omega}^{(i)}_{\text{pre}}\to 0$, then it has $\mathbb{E}[\mathbf{\xi}]\to 0$ and $\text{Var}[\mathbf{\xi}]\to 0$, so the $\mathbf{h}^{(i)}_{\text{pre}}$ have negligible influence on prediction logits. Even if the ideal scenario is hard to achieve, the classifier can still produce more accurate predictions as long as the representations adapted by reliable plugins are assigned higher weights, while those adapted by non-target plugins are appropriately suppressed. After enabling the weighting unit, $\mathcal{L}_{\text{IA}}$ is incorporated into the LoRA training, and $\mathcal{L}_{\text{LoRA}}$ in Eq. (8) is extended as follows:
$\mathcal{L}_{\text{LoRA}}=\mathcal{L}_{\text{CE}}+\mathcal{L}_{\text{aux}}+\mathcal{L}_{\text{IA}}.$ (12)
Table 1: Last performance $A_{T}$ and average performance $\overline{A}$ comparison. ‘-DLC’ indicates the method using the DLC framework. ‘LoRA #P’ denotes the total parameter count of the LoRA plugins, and ‘Ext #P’ refers to the total parameter count of the feature extractor.
@@ 5 Experiments
This section conducts extensive experiments. Sec. 5.2 evaluates the validity of DLC on replay and distillation-based baselines. Sec. 5.3 evaluates the compatibility of DLC on other data augmentation and plug-and-play loss function for CIL. Sec. 5.4 investigates the performance of baselines under memory-aligned case. An ablation study is conducted in 5.5 to analyze the contribution of the weighting unit.
@@ 5.1 Experimental Setup
Datasets and Split. We chose three benchmark datasets commonly used in CIL, including CIFAR-100 [krizhevsky2009learning], ImageNet-100 [deng2009imagenet], and Tiny-ImageNet-200 [vinyals2016matching], where the number following the dataset is the total number of classes included. Following the ‘B/Base-m, Inc-n’ rule proposed by [de2021continual], we slice each of the above three datasets into CIFAR-100 B10 Inc10, CIFAR-100 B50 Inc10, CIFAR-100 B5 INC5, ImageNet-100 B50 Inc10, and Tiny-ImageNet B40 Inc40 for the experiment, where $m$ is the number of classes in the first incremental task, and $n$ represents that of every remaining stage. For a fair comparison, we ensure that the training and test datasets are identical for all methods. Detailed information about the dataset is placed in the supplementary materials.
@@
Implementation Details. All methods are implemented in PyTorch, with the baseline methods referencing the PyCIL [zhou2023pycil] and Gao et al. [gao2025maintaining]. We employ the lightweight ResNet-32 tailored for CIFAR datasets on CIFAR-100, while utilizing ResNet-18 for all other benchmark datasets. The optimizer employs SGD with a fixed learning rate of 0.1, incorporating a multi-step schedule. The batch size is configured at 128 for CIFAR-100 and 256 for all other datasets. The number of training epochs was set consistent with the baseline method. All the methods are run using the same random number seed.
@@
Evaluation Metrics. Following the benchmark protocol settings [rebuffi2017icarl], we use $A_{b}$ to denote the $b$-stage accuracy on the test set that includes all known classes after trained with $D_{1},D_{2},\cdots,D_{b}$, $\overline{A}=\frac{1}{T}\sum_{b=1}^{T}A_{b}$ is average accuracy over $T$ tasks, and $A_{T}$ is the last accuracy on the test set that includes all learned tasks. We use $\overline{A}$ and $A_{T}$ to evaluate the model’s performance, which reflects the model’s dynamic learning capacity and generalization ability, respectively.
5.2 Evaluation on Non-Expansion Baselines
@@
In this section, we evaluate the performance improvement brought by the proposed DLC framework when integrated with established distillation-based baselines. We select four representative methods—LwF [li2017learning], iCaRL [rebuffi2017icarl], WA [zhao2020maintaining], and BiC [wu2019large] —where LwF (without a replay buffer) is used to assess DLC’s enhancement effect in a pure distillation setting. Additionally, a simple Replay method is included to examine DLC’s contribution in scenarios relying solely on the herding rehearsal without distillation. We report the $\overline{A}$ and $A_{T}$ of these established methods with and without DLC across three datasets under varying task configurations. In all experiments, the replay buffer size is fixed at 2,000 exemplars. DLC expands only one LoRA plugin per task, applied to the last convolutional layer of the feature extractor. We also report the total parameters required by the LoRA plugins and the size of the entire feature extractor for comparison. The results are summarized in Table 1.
@@
The results show that DLC significantly improves all classical replay or distillation-based methods with only one LoRA plugin per task. Moreover, the parameter overhead introduced by DLC is negligible. DLC achieves up to 8% accuracy improvement while expanding only 4.4% to 22% parameters equivalent to the feature extractor, highlighting its practical value in CIL. Fig. 3 shows the confusion matrix heatmap of iCaRL and iCaRL-DLC as an example. DLC increases the thermal intensity along the diagonal for most classes, demonstrating that it could enhance the model’s accuracy across most tasks.
@@
5.3 Compatibility with other Plug-and-Play Enhancements
In addition to DLC, other plug-and-play strategies have been proposed to enhance CIL methods. For example, the data augmentation technique AutoAugment [cubuk2018autoaugment] has been employed in the CIFAR-100 dataset by [wang2022foster, zheng2025task]. Meanwhile, two additional training losses $\mathcal{L}_{\text{inter}}$ and $\mathcal{L}_{\text{intra}}$ specifically designed for logit-based distillation methods have been introduced in [gao2025maintaining]. This subsection evaluates the compatibility of DLC with such enhance strategies. Specifically, we compare the accuracy of methods using AutoAugment and the loss $\mathcal{L}_{\text{inter-intra}}$ with and without DLC. As shown in the Fig. 4, DLC consistently brings significant improvements, demonstrating its compatibility with other plug-and-play enhancements.
@@ 5.4 Comparison under Fixed Memory Budget
Prior research [zhou2024class] has demonstrated that model-level expansion-based methods achieve state-of-the-art performance across all benchmarks. However, these approaches overlook the substantial memory overhead induced by dynamic architecture growth. Such methods implicitly introduce additional memory budgets, i.e., a model buffer for retaining parameters of the expanded model, which grants them a considerable and unaccounted advantage over other methods that do not store expansion models or adopt more lightweight expansion strategies.
Following the protocol established in [zhou2022model], this section presents a memory-aligned comparison of CIL baselines without pre-training. Under a fixed 23.5 MB memory budget, the memory not consumed by a method’s parameters is allocated to replay exemplars. To balance this, methods with fewer parameters store more exemplars, and vice versa. This ensures an equitable evaluation. In addition to the baselines used in Section 5.2, we add several state-of-the-art approaches for comprehensive comparison, including FOSTER [wang2022foster], DER [yan2021dynamically], MEMO [zhou2022model], and TagFex [zheng2025task]. Among them, DER, MEMO, and TagFex have incorporated the weight alignment from WA in their implementations. As summarized in Table 2, deploying DLC to the distillation-based method WA under the memory-aligned setting yields the highest classification accuracy.
@@
WA-DLC 23.5 0.72 7000 75.02 64.79
5.5 Ablation Study on the Weighting Unit
This subsection evaluates the impact of the proposed weighting unit on model performance. We compare the final accuracy of DLC-enhanced iCaRL, WA, and BiC with and without the weighting unit, as shown in Fig. 5. Results indicate that the proposed weighting unit contributes a certain accuracy gain. Although this improvement is relatively modest compared to the gains achieved by the LoRA plugin, the weighting unit could effectively direct the model’s attention to representations adapted by the pivotal LoRA plugin or key features of the sample, thereby enhancing their discriminative power. Specifically, we visualize the activation patterns of the last convolutional layer with the target task LoRA plugin using Grad-CAM [selvaraju2017grad], as illustrated in Fig. 6. For a “road” sample, despite both models correctly classifying it, the model equipped with the weighting unit effectively focuses on the road surface. Regarding a “pear” sample, when the weighting unit is ablated, the activation patterns remain inactive, indicating that representations adapted by the target LoRA contribute negligibly to classification. In contrast, with the weighting unit, the model not only correctly utilizes residuals from the target plugin but also focuses on discriminative regions, such as the cross‑section of the pear to improve the accuracy.
Figure 6: Grad-CAM visualizations for the last convolutional layer integrated with the target-task LoRA plugin. The model is iCaRL-DLC trained on the CIFAR-100 B10 Inc10. ‘Original’ denotes the input image, while ‘W/O weight’ and ‘Weighting’ represent the results without and with the weighting unit.

### v2

@@ 4.2 LoRA-Style Residual Plugin
where $\mathbf{A}\in\mathbb{R}^{r\times C_{\text{in}}\times K\times K}$ projects the feature to a low-rank space. This is followed by $\mathbf{B}$:
@@
where $\mathbf{B}$ $\in\mathbb{R}^{C{\text{out}}\times r\times 1\times 1}$ restores the channel dimensionality, and $\boldsymbol{h}_{\text{lora}}$ is added as a residual to the layer output.
In summary, the forward pass of a convolutional layer equipped with the LoRA-style residual plugin is:
$\boldsymbol{h}^{\prime}=W_{\text{conv}}\ast\boldsymbol{x}+\frac{\alpha}{r}(\mathbf{B}\ast(\mathbf{A}\ast\boldsymbol{x})),$ (9)
where $\alpha$ is a scaling hyperparameter, and $r$ is the rank.
@@
Compared to the model expansion, instantiating task-specific plugins with a LoRA-style parameterization can bring significant parameter advantages. Specifically, for a convolutional layer ($C_{in}$, $C_{out}$, $K\times K$), the space complexity of its parameters is $\Theta(C_{in}C_{out}K^{2})$, while the plugin only occupies $\Theta(r(C_{in}K^{2}+C_{out}))$. For expanding one layer of CNN over $T$ tasks, considering that in a CNN in CIL, $C_{out}$ is generally greater than or equal to $C_{in}$, we set $C_{in}=C_{out}=C$, and $K$ can be regarded as a constant (usually 3 or 7). Since $r$ is low-rank, it can also be regarded as a constant (generally set to 8 or 16). Then, using LoRA-style plugin instantiation reduces the parameter-space complexity of expanding a single CNN layer from $\Theta(TC^{2})$ to $\Theta(TC)$. Considering that $C$ is generally between 64 and 512, reducing its order of magnitude from quadratic to linear can substantially reduce the number of parameters.
@@ 4.3 Dynamic Weighting Unit
For a sample from task $t$, all the plugins trained on other tasks are considered non-target plugins. Given that the replay buffer contains only a minimal fraction of the original data for each task, these non-target plugins are not trained on the full data from task $t$. In particular, plugins trained on tasks 1 to t1t-1 never encounter samples from the task $t$. Consequently, the residuals extracted by these plugins could be largely irrelevant to the task $t$, thus introducing uninformative features or even detrimental noise to the classifier.
@@
where the $\mathbf{\omega}_{\text{ideal}}$ is the ideal weight output, it has $\mathbf{\omega}^{(i)}_{\text{ideal-pos}}=1$ and $\mathbf{\omega}^{(i)}_{\text{ideal-pre}}=0$. In the ideal training case, the model has $\mathbf{\omega}^{(i)}_{\text{pos}}\to 1$ and $\mathbf{\omega}^{(i)}_{\text{pre}}\to 0$, then it has $\mathbb{E}[\mathbf{\xi}]\to 0$ and $\text{Var}[\mathbf{\xi}]\to 0$, so the $\mathbf{h}^{(i)}_{\text{pre}}$ have negligible influence on prediction logits. Even if the ideal scenario is hard to achieve, the classifier can still produce more accurate predictions as long as the representations adapted by reliable plugins are assigned higher weights, while those adapted by non-target plugins are appropriately suppressed. After enabling the weighting unit, $\mathcal{L}_{\text{IA}}$ is incorporated into the plugin training, and $\mathcal{L}_{\text{plg}}$ in Eq. (5) is extended as follows:
$\mathcal{L}_{\text{plg}}=\mathcal{L}_{\text{CE}}+\mathcal{L}_{\text{aux}}+\mathcal{L}_{\text{IA}}.$ (12)
Table 1: Last performance $A_{T}$ and average performance $\overline{A}$ comparison. ‘-DLC’ indicates the method using the DLC framework. ‘Extra/Feat #P’ denotes the total number of extra parameters required for deploying DLC and the feature extractor, respectively.
@@ 5 Experiments
This section conducts extensive experiments. Sec. 5.2 evaluates the validity of DLC on distillation-based baselines. Sec. 5.3 investigates the performance of SOTA CIL methods under the memory-aligned case. Sec. 5.4 evaluates the compatibility of DLC with other data augmentation and plug-and-play loss functions for CIL. Sec. 5.5 analyzes the additional inference overhead introduced by DLC in terms of FLOPs and latency.
@@ 5.1 Experimental Setup
Datasets and Split. We chose three benchmark datasets commonly used in CIL, including CIFAR-100 [15], ImageNet-100 [9], and Tiny-ImageNet-200 [28], where the number following the dataset is the total number of classes included. Following the ‘B/Base-m, Inc-n’ rule proposed by [7], we slice each of the above three datasets into CIFAR-100 B10 Inc10, CIFAR-100 B50 Inc10, CIFAR-100 B5 Inc5, ImageNet-100 B50 Inc10, and Tiny-ImageNet B40 Inc40 for the experiment, where $m$ is the number of classes in the first incremental task, and $n$ represents that of every remaining stage. For a fair comparison, we ensure that the training and test datasets are identical for all methods. Detailed information about the dataset is placed in the Appendix section.
@@
Implementation Details. All methods are implemented in PyTorch, with the baseline methods referencing the PyCIL [41], and Gao et al. [10]. We employ the lightweight ResNet-32 tailored for CIFAR datasets on CIFAR-100, while utilizing ResNet-18 for all other benchmark datasets. The optimizer employs SGD with a fixed learning rate of 0.1, incorporating a multi-step schedule. The detailed training hyperparameters, like the optimizer and batch size, largely follow the default settings in PyCIL. All the methods are run using the same random number seed. To eliminate potential confounding effects from increased training epochs, all baselines are run under both the PyCIL-recommended epoch setting and the total epoch budget of DLC’s two-stage training, and the maximum accuracy is reported.
@@
Evaluation Metrics. Following the benchmark protocol settings [23], we use $A_{b}$ to denote the $b$-stage accuracy on the test set that includes all known classes after trained with $D_{1},D_{2},\cdots,D_{b}$, $\overline{A}=\frac{1}{T}\sum_{b=1}^{T}A_{b}$ is average accuracy over $T$ tasks, and $A_{T}$ is the last accuracy on the test set that includes all learned tasks. We use $\overline{A}$ and $A_{T}$ to evaluate the model’s performance, which reflects the model’s dynamic learning capacity and generalization ability, respectively.
5.2 Effectiveness on Distillation-Based CIL
@@
In this section, we evaluate the performance improvement brought by the proposed DLC framework across different distillation-based CIL objectives, including iCaRL [23], WA [38], BiC [31], DAKD [12], and MKD [20]. Additionally, a simple herding-based Replay method [23] is included to examine DLC’s contribution in scenarios relying solely on the rehearsal without distillation. We report the $\overline{A}$ and $A_{T}$ of these established methods with and without DLC across three datasets under varying task configurations. In all experiments, the replay buffer size is fixed at 2,000 exemplars. DLC expands only one LoRA plugin per task, applied to the last convolutional layer of the feature extractor. We also report the total parameters required by DLC and the size of the entire feature extractor for comparison. The results are summarized in Table 1.
@@
The results show that DLC significantly improves all classical replay or distillation-based methods with only one LoRA plugin per task. Moreover, the parameter overhead introduced by DLC is negligible. DLC achieves up to 8% accuracy improvement while expanding only 4.8% to 32% parameters equivalent to the feature extractor, highlighting its practical value in CIL. Fig. 3 shows the confusion matrix heatmap of iCaRL and iCaRL-DLC as an example. DLC increases the thermal intensity along the diagonal for most classes, demonstrating that it could enhance the model’s accuracy across most tasks.
In addition, we compare the final accuracy of DLC-enhanced iCaRL, WA, and BiC with and without the weighting unit, as shown in Fig. 4. Results indicate that the proposed weighting unit contributes a certain accuracy gain.
@@ 5.3 Comparison under Fixed Memory Budget
WA-DLC 23.5 0.54 7327 75.02 64.79
Prior research [42] has demonstrated that model-level expansion-based methods achieve state-of-the-art performance across all standard CIL benchmarks. However, these approaches overlook the substantial memory overhead induced by dynamic architecture growth. Such methods implicitly introduce additional memory budgets, i.e., a model buffer for retaining parameters of the expanded model, which grants them a considerable and unaccounted advantage over other methods that do not store expansion models or adopt more lightweight expansion strategies.
Following the protocol established in [43], this section presents a memory-aligned comparison of CIL baselines. Under a fixed 23.5 MB memory budget, the memory not consumed by a method’s parameters is allocated to replay exemplars. To balance this, methods with fewer parameters store more exemplars, and vice versa. This ensures an equitable evaluation. In addition to the baselines used in Section 5.2, we add several state-of-the-art approaches for comprehensive comparison, including FOSTER [29], DER [33], MEMO [43], and TagFex [39]. Among them, DER, MEMO, and TagFex have incorporated the weight alignment from WA in their implementations. As summarized in Table 2, deploying DLC to the distillation-based method WA under the memory-aligned setting yields the highest classification accuracy.
5.4 Compatibility with other Plug-and-Play Enhancements
In addition to DLC, other plug-and-play strategies have been proposed to enhance CIL methods. For example, the data augmentation technique AutoAugment [6] has been employed in the CIFAR-100 dataset by [29, 39]. Meanwhile, two additional training losses $\mathcal{L}_{\text{inter}}$ and $\mathcal{L}_{\text{intra}}$ specifically designed for logit-based distillation methods have been introduced in [10]. This subsection evaluates the compatibility of DLC with such enhanced strategies. Specifically, we compare the accuracy of methods using AutoAugment and the loss $\mathcal{L}_{\text{inter-intra}}$ with and without DLC. For experiments with $\mathcal{L}_{\text{inter-intra}}$, we follow the original implementation and use ResNet-18 as the backbone. As shown in the Fig. 5, DLC consistently brings significant improvements, demonstrating its compatibility with other plug-and-play enhancements.
5.5 Additional Computational Overhead from DLC
This section investigates the additional inference cost introduced by deploying DLC. Using iCaRL as a representative baseline, we evaluate two deployment variants: iCaRL equipped with one LoRA plugin per task, and iCaRL equipped with both the LoRA plugins and the weighting unit. Concretely, on ImageNet-100, we report the total FLOPs and the average inference latency per sample (ms) measured on the same test set. The results are shown in Fig. 6. Overall, neither FLOPs nor latency increases substantially as the number of tasks grows. Instead, both grow mildly as the task sequence extends. Even under long sequences such as B5 Inc5, the increases in FLOPs and inference time are only about 4%. These results indicate that DLC provides a lightweight and computation-friendly parameter extension for enhancing distillation-based CIL.
Figure 6: Comparison of inference FLOPs and per-sample inference latency with and w/o DLC under different task settings.